# PFP/MM: A Hybrid Approach Combining a Universal Neural Network Potential with Classical Force Fields for Large-Scale Reactive Simulations

### 17 Mar 2026 v1

@@ 3.3 P450-Cpd-I-Mediated Hydroxylation Reaction
Figure 4: (a) Structure of the P450-Cpd-I enzyme-ligand model used in the PFP/MM simulation. The PFP region is shown in color, while the MM region is shown in gray. (b) Hydroxylation reaction mechanism of P450-Cpd-I proposed in ref. Wang2021-sb. (c) Collective variables used in the metadynamics and umbrella sampling simulations. (d) Free energy surface obtained from umbrella sampling and the minimum free energy path. (e) Free energy profile along the reaction coordinate.
As an example that leverages the universality of PFP, we studied a substrate hydroxylation reaction catalyzed by the metalloenzyme P450 Compound I (Cpd-I) shown in Figure 4a. The proposed mechanism (Figure 4b) involves hydrogen atom abstraction (HAA) by the oxo ligand of Cpd-I, followed by rebound of the hydroxyl group to complete hydroxylation.Wang2021-sb, Prier2017-ps, Ren2016-id The PFP region comprised the substrate, the porphyrin, and the cysteine ligand coordinating the Fe center, whereas the remainder of the protein and the aqueous solvent were treated in the MM region. The covalent boundary was handled using the link-atom approach (see the Supporting Information).
We first performed pre-metadynamics to obtain a qualitative free-energy surface and representative structures (i), (ii), and (iii) (see Figure. S2 in Supporting Information). As collective variables, we used $d_{\mathrm{OHC}}=d_{\mathrm{CH}}-d_{\mathrm{OH}}$, which expresses whether the hydrogen atom is closer to the substrate C or to the oxo O, and $d_{\mathrm{COFe}}=d_{\mathrm{FeO}}-d_{\mathrm{CO}}$, which distinguishes O bound to Fe from O inserted into the substrate C (Figure 4c). To obtain a more quantitative landscape, we carried out umbrella sampling over 147 windows around the reactive region identified by metadynamics. Each window was equilibrated for $100\text{\,}\mathrm{ps}$ and sampled for $100\text{\,}\mathrm{ps}$, for a total of $29.4\text{\,}\mathrm{ns}$. The FES was reconstructed using MBAR,Shirts2008-tn and the minimum free-energy path was estimated using the string method (Figure 4d). The resulting FES indicates that HAA is the dominant free-energy barrier. Along the minimum free-energy path, the system proceeds directly to (iii); even when the trajectory passes through (ii), the hydroxyl group rapidly migrates to yield (iii) without an additional barrier. This behavior is consistent with previous QM/MM PES scans.Wang2021-sb Figure 4e shows the free-energy profile along the reaction coordinate, from which we estimate an activation free energy of $\sim$12\text{\,}\mathrm{k}\mathrm{c}\mathrm{a}\mathrm{l}\mathrm{/}\mathrm{m}\mathrm{o}\mathrm{l}$\$ for HAA and an overall free-energy difference of $\sim$48\text{\,}\mathrm{k}\mathrm{c}\mathrm{a}\mathrm{l}\mathrm{/}\mathrm{m}\mathrm{o}\mathrm{l}$\$ between the initial and final states.
@@ 4 Conclusion
We presented PFP/MM, a hybrid approach that combines the universal MLIP PreFerred Potential (PFP) with classical molecular mechanics (MM) by treating chemically active regions with PFP and the remaining majority of atoms with a fast MM description. This domain decomposition enables a favorable balance between accuracy and computational efficiency while retaining the broad transferability of a uMLIP. Benchmarks on alanine dipeptide in water demonstrated substantial speedups over PFP-only (17.8-fold on V100 and 56.5-fold on MN-Core 2 for a 7,975-atom system) and confirmed that nanosecond-per-day throughput remains feasible even at the million-atom scale. In one-day well-tempered metadynamics, PFP/MM achieved nanosecond-scale sampling and produced a stable Ramachandran free-energy surface reproducing the reported conformational basins, whereas PFP-only suffered from insufficient sampling. Recognizing that mechanical embedding limits the ability of the MM environment to polarize the PFP region, we showed that solvent polarity and hydrogen bonding effects can be incorporated by explicitly including nearby solvent molecules in the PFP region with FIRES. For an intramolecular nucleophilic addition, this strategy reproduced solvent-dependent stabilization of a zwitterion-like cyclized state in water relative to vacuum. Finally, we applied PFP/MM to a metalloenzyme reaction (P450-Cpd-I-mediated hydroxylation) and obtained a free-energy landscape consistent with the accepted mechanism, in which hydrogen atom abstraction constitutes the primary barrier and hydroxyl rebound proceeds without an additional barrier. Overall, PFP/MM provides a practical framework for large-scale, long-time reactive simulations that would be difficult to access with uMLIPs alone under realistic computational budgets.

### 24 Mar 2026 v2

@@ 3.3 P450-Cpd-I-Mediated Hydroxylation Reaction
Figure 4: (a) Structure of the P450-Cpd-I enzyme-ligand model used in the PFP/$M$ simulation. The PFP region is shown in color, while the $M$ region is shown in gray. (b) Hydroxylation reaction mechanism of P450-Cpd-I proposed in ref. Wang2021-sb. (c) Collective variables used in the metadynamics and umbrella sampling simulations. (d) Free energy surface obtained from umbrella sampling and the minimum free energy path. (e) Free energy profile along the reaction coordinate.
As an example that leverages the universality of PFP, we studied a substrate hydroxylation reaction catalyzed by the metalloenzyme P450 Compound I (Cpd-I) shown in Figure 4a. The proposed mechanism (Figure 4b) involves hydrogen atom abstraction (HAA) by the oxo ligand of Cpd-I, followed by rebound of the hydroxyl group to complete hydroxylation.Wang2021-sb, Prier2017-ps, Ren2016-id The PFP region comprised the substrate, the porphyrin, and the cysteine ligand coordinating the Fe center, whereas the remainder of the protein and the aqueous solvent were treated in the $M$ region. The covalent boundary was handled using the link-atom approach (see the Supporting Information).
We first performed pre-metadynamics to obtain a qualitative free-energy surface and representative structures (i), (ii), and (iii) (see Figure. S2 in Supporting Information). As collective variables, we used $d_{\mathrm{OHC}}=d_{\mathrm{CH}}-d_{\mathrm{OH}}$, which expresses whether the hydrogen atom is closer to the substrate C or to the oxo O, and $d_{\mathrm{COFe}}=d_{\mathrm{FeO}}-d_{\mathrm{CO}}$, which distinguishes O bound to Fe from O inserted into the substrate C (Figure 4c). To obtain a more quantitative landscape, we carried out umbrella sampling over 147 windows around the reactive region identified by metadynamics. Each window was equilibrated for $100\text{\,}\mathrm{ps}$ and sampled for $100\text{\,}\mathrm{ps}$, for a total of $29.4\text{\,}\mathrm{ns}$. The FES was reconstructed using MBAR,Shirts2008-tn and the minimum free-energy path was estimated using the string method (Figure 4d). The resulting FES indicates that HAA is the dominant free-energy barrier. Along the minimum free-energy path, the system proceeds directly to (iii); even when the trajectory passes through (ii), the hydroxyl group rapidly migrates to yield (iii) without an additional barrier. This behavior is consistent with previous QM/$M$ PES scans.Wang2021-sb Figure 4e shows the free-energy profile along the reaction coordinate, from which we estimate an activation free energy of $\sim$12\text{\,}\mathrm{k}\mathrm{c}\mathrm{a}\mathrm{l}\mathrm{/}\mathrm{m}\mathrm{o}\mathrm{l}$\$ for HAA and an overall free-energy difference of $\sim$48\text{\,}\mathrm{k}\mathrm{c}\mathrm{a}\mathrm{l}\mathrm{/}\mathrm{m}\mathrm{o}\mathrm{l}$\$ between the initial and final states.
@@ 4 Conclusion
We presented PFP/$M$, a hybrid approach that combines the universal MLIP PreFerred Potential (PFP) with classical molecular mechanics ($M$) by treating chemically active regions with PFP and the remaining majority of atoms with a fast $M$ description. This domain decomposition enables a favorable balance between accuracy and computational efficiency while retaining the broad transferability of a uMLIP. Benchmarks on alanine dipeptide in water demonstrated substantial speedups over PFP-only (17.8-fold on V100 and 56.5-fold on MN-Core 2 for a 7,975-atom system) and confirmed that nanosecond-per-day throughput remains feasible even at the million-atom scale. In one-day well-tempered metadynamics, PFP/$M$ achieved nanosecond-scale sampling and produced a stable Ramachandran free-energy surface reproducing the reported conformational basins, whereas PFP-only suffered from insufficient sampling. Recognizing that mechanical embedding limits the ability of the $M$ environment to polarize the PFP region, we showed that solvent polarity and hydrogen bonding effects can be incorporated by explicitly including nearby solvent molecules in the PFP region with FIRES. For an intramolecular nucleophilic addition, this strategy reproduced solvent-dependent stabilization of a zwitterion-like cyclized state in water relative to vacuum. Finally, we applied PFP/$M$ to a metalloenzyme reaction (P450-Cpd-I-mediated hydroxylation) and obtained a free-energy landscape consistent with the accepted mechanism, in which hydrogen atom abstraction constitutes the primary barrier and hydroxyl rebound proceeds without an additional barrier. Overall, PFP/$M$ provides a practical framework for large-scale, long-time reactive simulations that would be difficult to access with uMLIPs alone under realistic computational budgets.